OpenStudy (anonymous):

Multiply.Use the distributive property. A) (3x+2y-1)(xy)

OpenStudy (zehanz):

Distributive property: a(b+c)=ab+ac, or to make it look more like your problem: (a+b+c)d = ad+bd+cd. It's called "distributive, because the multiplication with d distributes over every term between the brackets.

OpenStudy (zehanz):

@jenny508: Let me help you a bit further: (3x+2y-1)(xy) = 3x*xy + ...